### v1

# An analytical model for the remote epitaxial potential

Abstract
We propose an analytical model for the remote bonding potential of the substrate that permeates through graphene during remote epitaxy. Our model, based on a Morse interatomic potential, includes the attenuation due to (1) the increased separation between film and substrate and (2) free carrier screening from graphene. Compared with previous slab density functional theory calculations, which use the electrostatic potential as a proxy for bonding, our analytical model provides a more direct description of bonding, explicitly includes screening (which is often ignored), and is based on simple, interpretable, and well benchmarked parameters. We show that the magnitude of for typical semiconductor and oxide substrates is few meV, similar to the van der Waals potential of graphene. This suggests interference between the graphene and remote substrate potentials, plus interfacial reconstructions, must be considered when interpreting experiments on remote epitaxy. We use our model to interpret previous experiments from the remote epitaxy and related literature, highlighting connections to moire epitaxy and to the ordering of graphite intercalation compounds. Our model also points to tests, based on tunable screening and spatial extent of the substrate potential, that may increase the strength of the remote potential towards the more idealized picture.
Bonding is typically considered a very short-range interaction ( 3 Å). Atomic and molecular orbitals (wavefunctions) decay exponentially with distance, and thus nearest neighbor interactions tend to dominate. Remote epitaxy challenges this paradigm. Here it is argued that the lattice potential of a crystalline substrate can permeate through graphene with sufficient strength to template epitaxy of a film on top of a graphene-covered substrate [1, 2]. In this picture, the graphene is nearly transparent [3]: the film adopts epitaxial registry to the buried substrate rather than graphene, and the lattice potential of the graphene itself is assumed to contribute minimally. The graphene spacer merely attenuates the film-substrate interaction strength. This concept is quite useful: due to the weaker film-substrate interactions, films grown on graphene can be exfoliated to produce flexible, stackable, and strain engineered [4, 5] membranes. The weakened interactions may also enable lattice mismatched heteroepitaxy of films with decreased dislocation density due to the “slippery” graphene interface.
The current theoretical understanding of remote epitaxy is based on the density functional theory (DFT)-computed electrostatic potential decay with distance[2, 6, 7, 8], which typically ignore interfacial reconstructions and assumes all the atoms are near their bulk like positions (presumably to minimize computational cost). These calculations suggest that substrates with greater bond polarity produce larger electrostatic fluctuations above graphene, and thus can support remote epitaxy through greater number of graphene layers [2].
However, there are several challenges to this understanding. First, it is not clear that a purely electrostatic interaction can explain remote epitaxy, since the electrostatic potential describes the electronic workfunction rather than bonding (i.e., the energy of pulling an electron to infinity from that location). Electrostatic potentials are thus completely agnostic to the chemical match between the film and substrate, which has been shown to impact remote epitaxy efficacy [9]. Second, most DFT treatments ignore free carrier screening from graphene, ignore the lattice potential of graphene, and ignore possible reconstructions. Finally, DFT calculations can be a black box, making it difficult to interpret the contributions from various terms without additional tests. A simple analytical model is needed in order to build a physical understanding of the various terms that contribute to the strength of the remote potential.
Here we develop an analytical model for the remote bonding potential permeation through graphene. Our model is based on an empirical Morse interatomic potential [10], which includes contributions from both covalent bonding and electrostatic interactions [11]. This simple two body potential captures the essential of physics and employs easy to understand parameters that are well benchmarked by experiments, density functional theory, and molecular dynamics simulations [12]. Compared to previous approaches that used the electrostatic potential, our approach (1) uses a potential that includes both covalent bonding and electrostatic interactions [11], (2) explicitly includes free carrier screening of graphene, which is often ignored [6, 1, 8], (3) avoids the use of an adjustable scale factor that has been invoked to make the electrostatic potential behave more like a bonding potential [2], and (4) avoids the use of complicated background subtraction and data cropping [2]. We show that graphene is not as transparent as often assumed: the remote potential of the substrate through graphene has an order of magnitude few meV or smaller, similar to the the vdW interaction strength of graphene itself. Our model predicts that the spatial extent of the bond potential, rather than bond ionicity, is the primary parameter that controls the strength of the remote potential.
Our model treats graphene in ideal limit as a homogeneous electron gas. The key elements are (1) increased separation between film and substrate due to insertion of the graphene and (2) attenuation due to free carrier screening. We start from a Morse potential for the substrate
where the first term describes the short-range Pauli repulsion and the second term is the long range attraction. Here is the bond strength, is the width of the potential well, and is the equilibrium bond distance. The potential for a bare GaAs substrate without a remote spacer layer is shown in Fig. 1(a) (black curve). This potential describes the pair bonding interaction between an As atom and a Ga atom, and thus is appropriate for homoepitaxy of a GaAs film on a GaAs substrate. We use parameters eV and Å-1, based on optical spectroscopy measurements of the bond strength and oscillation frequency of GaAs [13], and refined by ab initio calculations in a many-body form to reproduce the bulk properties [14, 15]. We let Å, corresponding to a generalized layer spacing between atomic planes.
We next model the insertion of an insulating 2D material spacer at distance from the substrate, in the absence of screening. This is analogous to inserting monolayer hBN, to produce a new minimum at increased distance from the substrate. Since the primary out of plane interactions for 2D materials are van der Waals, we model the 2D material with a Lennard-Jones potential
where we let meV and Å. The resulting potential , in the absence of screening, is shown in the blue curve. The inert spacer displays a new minimum of meV, which is only of the Ga-As bond energy. The new equilibrium film-substrate separation of Å is within the 5-6 Å separation typically observed by TEM of GaAs/graphene/GaAs [16]. Our insulating spacer calculation is consistent with previous slab DFT calculations that use charge density as a proxy for the remote potential permeation: for (001)-oriented GaAs slabs separated by a 7 Å vacuum gap (no graphene for screening), the computed charge density in the middle of the gap was of the charge density in the interior of the material [1].
We now include free carrier screening from metallic graphene. This screening is typically ignored by slab DFT calculations that either do not include or try to subtract out the contributions from graphene [1, 2, 6]. We model the graphene as a homogeneous electron gas with Thomas-Fermi wavenumber where is the carrier density, is the electron rest mass, and is the reduced Planck constant. The screened transmission coefficient through graphene is
where Å is the effective thickness of graphene and is the number of graphene layers. Figure 1(c) plots through monolayer and bilayer graphene, as a function of typical room temperature graphene carrier densities cm-2 (). We find that the screening is highly tunable with carrier density, and there is a significant screening-induced attenuation going from monolayer graphene to bilayer graphene (in addition to the attenuation from increased layer separation). For greater than the thickness of graphene, the total screened potential is
where the first term is the screened remote potential of the substrate and the second term is the Lennard-Jones (vdW) potential of graphene.
Fig. 1(b) plots for a range of typical graphene carrier densities. Here our model provides two testable predictions. First, the equilibrium film-substrate separation (minimum of ) shifts to larger with increased carrier density (screening). Thus the separation can serve as a proxy for the remote interaction strength. For very large graphene carrier density, or as approaches 0, converges to the Lennard-Jones potential of graphene . Electrostatic gating of the graphene could be used to test this prediction.
Second, the magnitude of is significantly attenuated by screening. For a carrier density of cm-2, the magnitude of the at the new minimum of Å is 15 meV. This magnitude is smaller than the GaAs covalent bond strength ( eV), smaller than typical surface diffusion barrier heights (100 meV), and comparable to the meV scale for van der Waals interactions (Fig. 1(c)). Note that the template for epitaxy is given by lateral fluctuations in the potential rather than the full magnitude of the total potential . This suggests that the lateral fluctuations are smaller than 15 meV.
Our calculations reveal that once screening is included, the remote potential of GaAs through monolayer graphene is not the dominant interaction. Rather, has similar magnitude as the meV scale van der Waals interaction of graphene itself, and hence the graphene potential cannot be ignored in remote epitaxy. Since this screening depends exponentially on the thickness of the metallic layer , the prospect for “remote” epitaxy through bilayer or even trilayer graphene is vanishingly weak, in contrast with previous experimental reports of remote epitaxy through several layers of graphene [2]. Note that most experiments cannot rule our pinhole nucleation as the mechanism rather than remote interactions, since contaminants and oxides at the transferred graphene/substrate interfaces typically create pinholes in the graphene when heated to film growth temperatures [17].
At first glance this may seem inconsistent with previous DFT slab calculations of the permeated potential, which suggested 15 meV fluctuations through graphene on GaAs substrates [2]. However, we emphasize that (1) the previous calculation was for the electrostatic potential [2], which is not the bonding potential, (2) the calculations employed selective cropping and replication and thus is not a reliable representation of the potential fluctuations [2], and (3) in many calculations the screening effects from graphene and the lattice potential of graphene itself are not explicitly included.
(eV) | (Å-1) | (meV) | |||
ZnO | 1.79 | 3.60 | 1.82 | 5.3 | |
GaN | 1.63 | 2.45 | 1.97 | 2.3 | |
GaAs | 0.37 | 2.10 | 1.52 | 7.6 | |
SiC | 0.65 | 4.36 | 1.70 | 9.3 | |
Si | 0 | 3.24 | 1.48 | 13 | |
Cu | 0 | 0.34 | 2.87 | 0.022 |
We now apply our model to other commonly used semiconductor and oxide substrates. Table 1 summarizes the Morse potential parameters for ZnO [18], GaN [19], GaAs [14, 15], SiC [20], Si [20], and Cu [21]. We first use these parameters to estimate the attenuation from the separation induced by the offset of an inert 2D materials spacer, compared to the attenuation from screening. For simplicity, we use the fact that for large the attractive term of the Morse potential (Eqn 1) dominates. We thus use . In Table 1 we compute at Å. We find that the separation induced attenuation for these substrates of to is slightly more significant than the screening induced for monolayer graphene at typical carrier densities cm-2 (Fig. 1(b)).
We next test the hypothesis from Ref. [2] that polar ionic bonding controls the strength of the remote potential. Table 1 summarizes the remote potential under the approximation
We plot this magnitude versus the Pauling electronegativity difference in Fig. 2, where we let . We find no clear trend of with . On the other hand, a clear trend emerges with the Morse potential width , where more spatially extended bond potentials (small ) lead to a larger remote potential at a distance of 6 Å from the substrate. This trend is clearly seen by the shape of the potential (Eqn 1), where the potential varies exponentially with but only linearly on the bond strength .
Interestingly, our model suggests that monoatomic Si should have the strongest remote potential among the substrates considered. This lies in apparent contradiction to the experiments in Ref. [2] reported that growth of Si on graphene/Si (001) resulting in a polycrystalline film, yet growth on substrates with more polar bonding enabled epitaxial films in more ionic LiF substrates that were covered with up to three layers of graphene. However, the impact of key synthesis parameters including substrate temperature, fluxes, type of precursor (e.g., atom, molecular cluster, metalorganic), and graphene surface preparation remain unclear.
Our model calls for a re-examination of Si and other substrates with small , coupled with a systematic understanding of how synthesis conditions and graphene surface preparation affect the resulting films and their crystalline orientations. The model suggests that the best substrates for remote epitaxy are those with spatially extended bond potentials (or spatially extended orbitals), rather than substrates with more ionic bonding character. Our calculations also suggest that tuning the screening via gating the graphene may be a powerful route to tune the strength of the remote potential.
We caution that our model should only be considered an order of magnitude estimate for the remote potential. We treat graphene as an idealized homogeneous electron gas, and only consider pair-wise atom-atom interactions between film and substrate. Our model does not consider the directional character of bonding, multi-atom interactions, or bonding between the film and graphene or graphene and the substrate. It also does not consider surface reconstructions. Still, it identifies an essential role for the decay with distance and for screening. These essential parameters place significant constraints on the magnitude of the remote bonding potential that are overlooked in the current atomistic calculations. Experimental techniques such as He atom scattering [22] and force-distance spectra by atomic force microscopy [23, 24, 25, 26], as a function of graphene gating, could directly measure the remote potential as a function of screening to test our predictions.
Given the relatively weak remote potential, it is worth considering other explanations for remote epitaxy, beyond the simple picture of transparent graphene. For the bulk graphite intercalations compounds, the intercalants form ordered superlattices that can be commensurate or incommensurate to the graphite lattice [27]. This intercalant ordering is thought to be governed by several factors, including (1) Coulomb repulsion between intercalants, since there is charge transfer between the intercalants (often alkali metals) and the graphite sheets, and (2) strain fields in the graphite sheets induced by bending around the intercalants [28]. These factors may also be relevant to remote epitaxy. The former charge transfer is not a remote interaction per se, but hints at a possible extended molecular orbital picture that includes atomic orbitals from the film and substrate, hybridized with the extended system of graphene. Such a description is not included in our model. The second factor, strain and buckling in graphite, suggests that surface reconstructions may be relevant to epitaxy on graphene-covered substrates. Indeed, strong buckling is observed in the “buffer” layer of graphene that forms on Si-face SiC (0001) [29]. This distorted graphite has a mix of and bonds. The out of plane dangling bonds are expected produce bonding potential fluctuations of order 100 meV to eV (since they are covalent bonds), larger than the few meV scale remote and vdW potentials. Indeed, recent experiments demonstrate the adatom sticking coefficient is higher on buffer graphene than on epitaxial graphene [30], implying higher reactivity due to these dangling bonds. These reconstructions and moire interference with the graphene are known to template the ordering of metal adatoms on graphene/Ir (111) [31], graphene/Rh (111) [32], graphene/Pt (111) [33], graphene/SiC (0001) [34]. Atomistic calculations that include these reconstructions as well as direct chemical interaction between the film and substrate are required to fully understand remote epitaxy of continuous films on graphene.
I Acknowledgments
Model development by JKK was primarily supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under award no. DE-SC0023958. Comparison of different substrates was supported by the Sandia University Partnerships Network (SUPN) program of Sandia National Laboratories.
QTC is supported by the by the Laboratory Directed Research and Development (LDRD) program at Sandia National Laboratories under project 233271. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan.